Please help with finding blocking probablity when;
-M/M/1/1 queuing model
-Generates a 1000 packet UDP stream with lambda=200 packets per second, each packet being 100 bytes.
-Mean value of the processing rate of packets (mu)=50 packets per second.
Answer #1

Hi,

I have provided the explaination so that it will be more clear. If you have any other question reach out to me in the comment section. Please like it if you find it useful.

(a). It is known that the mean number of packets in this system is given by ρ/ρ-1

where, ρ=λ/μ

λ is the rate at which packets arrive at the queue (in packets/second), and μ is the rate at which packets may be served by the queue (in packets/second).

Here, λ = 200 packets/sec and μ = 50 packets/sec

ρ=λ/μ = 200/50 = 4

ρ/ρ-1 = 4/(4-1) = 1.33

Packet size = 100 bytes

M/M/1/K Queue (single server queue with K-1 waiting positions) a = b k<K otherwise (Blocking or Loss Condition) H* = 4 ksk =

Here K = 1

p0 = (ρ-1)/(ρK+1 - 1) = (4-1)/((4)1+1 - 1) = 3/15 = 0.2

Blocking probability pk = p0 ρk = (0.2) * (4)1 = 0.8 = 80%